OpenStudy (anonymous):

p=2(l+w) solve for l

OpenStudy (luigi0210):

Distribute and solve\

OpenStudy (anonymous):

ok so p=2l + 2w

OpenStudy (luigi0210):

Yes now move it over and solve for l p-2w=2l p/2-w=l

OpenStudy (rajee_sam):

\[p = 2(l + w) \]\[\frac{ p }{ 2 } = l + w\] \[\frac{ p }{ 2 } - w = l \]

OpenStudy (anonymous):

Thanks :) Buddies